# An improved description of nuclear charge radii: Global trends beyond $N=28$ shell closure

### v1

@@ III.1 Shell quenching of nuclear charge radii at $N=28$
where $R(N,Z)$ is rms charge radius for a nucleus with neutron number $N$ and proton number $Z$. In Figs. 1 (c) and (d), the OES of charge radii in potassium and calcium isotopes is depicted. For potassium isotopes, the OES in charge radii can be reproduced well across $N=20$ shell closure. Extending to neutron-deficient regions, the obtained OES amplitudes for these three effective forces cannot reproduce the experimental data well. This can be due to the fact that the charge radii in neutron-deficient region are overestimated by NL3 and $*{}^{*}$ sets and underestimated by PK1 force. For neutron-rich region, the PK1 force gives the enlarged OES of charge radii respect to NL3 and $*{}^{*}$ sets. Along calcium isotopes, the OES behaviors of charge radii for ${}^{37-47}$a isotopes can be reproduced well by these forces. Across the $N=28$ filled shell, PK1 set gives the inverse OES amplitudes in charge radii with respect to NL3 and $*{}^{*}$ sets. Besides, it can be found that the OES behavior is weakened at the neutron number $N=28$. This is in accord with Ref. An et al. (2022b) where the weakened OES in charge radii can be regarded as signature of the emergence of neutron magic numbers.
@@
Figure 2: Same as Fig. 1, but for scandium (a), titanium (b), vanadium (c), chromium (d), manganese (e), iron (f), cobalt (g), nickel (h) and copper (i) isotopes. Experimental data are taken from Refs. Angeli and Marinova (2013); Li et al. (2021), while the data of Sc isotopes are taken from Ref. König et al. (2023b), the value of ${}^{54}$i is extracted from Ref. Pineda et al. (2021) and ${}^{58-68,70}$i are extracted from Ref. Malbrunot-Ettenauer et al. (2022).
@@
To facilitate the instructive understanding of charge radii on the isotopic chains of neighbouring elements, the systematic evolution of charge radii along $Z=21-29$ isotopic chains is plotted in Fig. 2. From this figure, one can find that the shell effect at $N=28$ is evidently observed. Across $N=28$, the abrupt changes are universally observed in this region, while the slope of change is lower for scandium ($Z=21$) and titanium ($Z=22$). In Fig 2 (a), these three forces give the inverted parabolic-like behaviors in charge radii between $N=20$ and $N=28$, while PK1 parameter set slightly underestimates the results. The latest study suggests that the charge radii of ${}^{41}$c isotope is surprisingly shrunk König et al. (2023b), then the abrupt increase is occurred beyond the neutron number $N=20$. In our calculations, the PK1 force can reproduce the charge radii of ${}^{40,41}$c isotopes, but the values of ${}^{42-46}$c isotopes are exaggeratedly underestimated. By contrast, these phenomena are mostly melt by the NL3 and $*{}^{*}$ forces. The NL3 and $*{}^{*}$ forces can cover the experimental uncertainties of ${}^{40,41}$c, but the values of ${}^{42,43}$c are underestimated. It is notably mentioned that our calculations can give the trend of kink, but the amplitude is heavily reduced. Besides, the last unpaired nucleon is just tackled by the blocking approximation, namely occupies the last single-particle level. This may also result in the underestimated results. The same scenarios are also encountered for titanium isotopes. As shown in Ref. Gangrsky et al. (2004), the data of the charge radii on the titanium isotopes show a continuous decrease from $N=22$ to $28$. The calculated results for ${}^{44-46}$i are slightly deviated from experimental data. But the trend of changes of charge radii are not changed, especially for ${}^{44-46}$i where the odd-even oscillation behaviors can be reproduced well.
@@
For V, Cr and Co isotopic chains, the experimental data can be reproduced well. However, more experimental data are urgently encouraged to support our results. In these regions the overall slope increases gradually with the increasing neutron numbers. Also the shell closure effect in charge radii is clearly pronounced at $N=28$ due to the rather small isospin dependence of spin-orbit interactions Sharma et al. (1995). Along manganese isotopes, results obtained by NL3, PK1 and $*{}^{*}$ forces can reproduce the experimental data across $N=28$ shell closure. However, there exists a slight deviation between experimental data and results obtained by PK1 force. For ${}^{50,51}$n and ${}^{52}$e, the calculated results are all reduced. Meanwhile, the data of ${}^{56-58}$e isotopes are also slightly underestimated.
@@
Recent studies on nuclear size are performed along nickel isotopic chain Pineda et al. (2021); Malbrunot-Ettenauer et al. (2022). In Fig. 2 (h), one can find that NL3, PK1 and $*{}^{*}$ forces can describe the general trend of changes of charge radii along nickel isotopic chain. Especially the slight odd-even oscillation behavior can be reproduced well against experimental data. In Ref. Malbrunot-Ettenauer et al. (2022), it is mentioned that the trend of changes of charge radii for ${}^{62-70}\mathrm{Ni}$ isotopes are overestimated by Fayans EDF model. The sophisticated Fayans energy density functional (EDF) can reproduce the staggering effects of charge radii for Ca isotopes Reinhard and Nazarewicz (2017). Furthermore, the inverted parabolic-like behavior between $N=20$ and $N=28$ can also be reproduced well. As demonstrated in Ref. Malbrunot-Ettenauer et al. (2022), the isovector component is excluded in pairing interactions. This maybe lead to the lack of an isovector component in its pairing part. In Ref. Kortelainen et al. (2022), the charge radii of even-even Ca-Zn nuclei are obtained by the $ab$ $initio$ coupled cluster theory and nuclear density functional theory. In which a universal pattern from neutron numbers $N=28$ to $N=40$ is exhibited, namely the change in charge radii shows weak dependence on the atomic number. Actually, our results are in accord with those.
@@
The pioneering work shows that the charge radii formula without considering the unpaired nucleons correlation give the overestimated OES of charge radii in copper isotopes An et al. (2022c). As shown in Fig. 2 (i), the improved formula can reduce the overestimated odd-even oscillations of charge radii. In Ref. de Groote et al. (2020), it is found that the charge radii along copper isotopic chain are also slightly overestimated by Fayans DFT model. Especially close to neutron-rich region, this overestimated OES behaviors are evidently represented. Therefore, the further performance of optimization-based machine learning method is addressed through Fayans EDF model Bollapragada et al. (2021).
III.2 Charge radii of $N=28$, $32$ and $34$ isotones
@@
The emergence of rapidly increasing in charge radii is commonly observed across the traditional neutron magic numbers $N=28$, $50$, $82$ and $126$ in the nuclear chart Kreim et al. (2014); Gorges et al. (2019); Goddard et al. (2013); Procter et al. (2012). These phenomena arise from the relative stable binding properties with respect to the adjacent nuclei. Meanwhile, the new magicity of $N=32$ and $34$ has also been identified by multi-aspects. Recent studies show that the isotopes with neutron number $N=32$ are proposed to exhibit localized magic behavior due to an observed sudden decrease in binding energy beyond $N=32$, such as in ${}^{51}$ ($Z=19$Rosenbusch et al. (2015), ${}^{52}$a ($Z=20$An et al. (2018); Wienholtz et al. (2013); Gallant et al. (2012), ${}^{53}$c ($Z=21$Xu et al. (2019) and ${}^{54}$i ($Z=22$Janssens et al. (2002); Leistenschneider et al. (2018). This new magicity is also identified from the high excitation energy of the first excited state Steppenbeck et al. (2015); Huck et al. (1985); Liddick et al. (2004); Prisciandaro et al. (2001) and the reduced $B(E2)$ transition probabilities Goldkuhle et al. (2019); Seidlitz et al. (2011). Especially for ${}^{52}$a, the doubly-magic nature is also confirmed by two-proton knockout reaction Gade et al. (2006). As demonstrated in Ref. Reiter et al. (2018), the signature of $N=32$ sub-shell closure is not identified across the vanadium isotopic chain ($Z=23$) and higher proton numbers.
@@
Figure 3: Charge radii of $N=28$ (open circle), $32$ (open triangle) and $34$ (open square) isotones as a function of proton number are plotted within NL3 (a), PK1 (b) and $*{}^{*}$ (c) forces. The results obtained by the BNN+D4 (d) Dong et al. (2022), HFB-21 (e) Goriely et al. (2010) and WS (f) Wang and Li (2013) models are also shown for comparison. The corresponding experimental data are taken from Refs. Angeli and Marinova (2013); Li et al. (2021) (solid marks).
@@
In Fig. 3, the charge radii of nuclei along $N=28$, $32$ and $34$ isotonic chains are plotted. It clearly shows that the charge radii monotonically increase with the increasing proton number. As the slope of these curves involving different proton numbers is much changed, i.e., the local variations can be found obviously. The charge radii almost changes linearly from K ($Z=19$) to Fe ($Z=26$) along $N=28$ isotonic chain. The shrinking trend occurs until to nickel ($Z=28$), and then the abrupt change occurs across $Z=28$. This result suggests that the shell closure effect of charge radii can also be found along $Z=28$ isotones. These calculations for the rms charge radii reveal the characteristic kink at the $Z=28$ shell closure in accordance with the corresponding experimental radii.
@@
Along $N=32$ isotonic chain, the abrupt changes of charge radii across nuclei with $Z=22$ are observed within NL3 (a), PK1 (b) and $*{}^{*}$ (c) effective forces. From V to Ga isotopes, the charge radii are changed linearly with the increasing proton number, but the results of ${}^{58}$e are slightly deviated. This can be seen from the Fig. 2 that charge radii of Fe isotopes are systematically underestimated by these three forces. From K to Ti isotopes, the increasing trend of charge radii are almost linear. Same scenario can also be found along $N=34$ isotonic chain, but the slope of changes of charge radii from $Z=22$ to $Z=23$ is smaller. As discussed above, magicity means the nuclei process relatively stable properties. This pronounced feature can also be remarkably reflected by nuclear size. As suggested in Ref. Reiter et al. (2018), the signature of $N=32$ sub-shell closure is not identified across $Z=23$ and higher proton numbers. This is consistent with our results where the charge radii keep linear increase after the abrupt increase at $Z=22$. This may provide a signature to verify the new magicity of nuclei with $N=32$. The same scenario can also be observed along $N=34$ isotonic chain, but the slope of changes of charge radii is smaller. Consequently, more accurate experimental data are urgently needed in proceeding future.
@@
The latest experimental data show that there has no sharp change in charge radii across $N=32$ Koszorús et al. (2021). Although the magicity of nuclei with $N=32$ has been performed, the shrinking phenomena are not observed along potassium Koszorús et al. (2019) and calcium Garcia Ruiz et al. (2016) isotopes. A more complete picture may shed light on the unexplained phenomena and reveal details of the proton-neutron interaction under the influence of the $N=32$ and $34$ sub-shell closure. The $N=34$ sub-shell closure has been so far performed in Ca Michimasa et al. (2018); Steppenbeck et al. (2013); Chen et al. (2019) and Ar Liu et al. (2019) isotopes. However, the slightly sharp transition in charge radii is also observed across $Z=22$. A quantitative theoretical interpretation of the well-known sub-shell effect in $N$ has been hampered by the fact that detailed calculations use somewhat inconsistent models in treating even-$N$ and odd-$N$ nuclei. The present exceptional case of a pronounced $Z$-dependence of the charge radii along $N=28$, $32$ and $34$ isotonic chains may provide the clue for a clear-cut phenomenological interpretation.
@@
Figure 4: Odd-even staggering of charge radii versus $N=28$ (a), $32$ (b) and $34$ (c) isotones as a function of proton number are plotted within NL3 (open diamond), PK1 (open square) and $*{}^{*}$ (solid triangle) forces. The corresponding experimental data are taken from Refs. Angeli and Marinova (2013); Li et al. (2021).
@@
Such a comparison should be helpful in providing an accessible understanding of nuclear charge radii and in making reliable predictions for those nuclei lacking an experimental value. Although the systematic evolution of the charge radii over a large range of proton numbers seems to be consistent with the presence of a noticeable sub-shell effects at $N=32$ and $34$, it misses the pronounced experimental data to identify this results. In order to further examine the universal trend of nuclear charge radii, it will be also instructive to gather similar data on the isotonic chains at neutron number $N=28$, $32$ and $34$. In Fig 3 (d), (e) and (f), the results obtained by Bayesian Neural Network (BNN+D4) Dong et al. (2022), HFB-21 Goriely et al. (2010) and WS Wang and Li (2013) are also shown for comparison.
@@
In Fig. 3 (d), the change of charge radii obtained by BNN+D4 is almost linear along $N=28$ isotonic chain. The abrupt increases are predicted along $N=32$ and $34$ isotonic chains at the proton numbers $Z=20$ and $28$, respectively. This can be understood clearly where the pronounced shell effects are emphasized in the calibrated protocol Dong et al. (2022). For HFB-21 model Goriely et al. (2010), the charge radii of $N=28$, $32$ and $34$ isotonic chains change linearly with the increasing proton numbers. While the slight shrinking occurs at $Z=27$, this maybe result from the unpaired proton. This means the strong shell effect is weakened in Skyrme functionals. As demonstrated in Refs. Tajima et al. (1993); Gorges et al. (2019), the calculations in non-relativistic density functional theories (DFT) based on conventional functionals are unable to reproduce the kink well at magic numbers. Such as Skyrme functionals utilizing SV-min force generally cannot reproduce the characteristic kink typically found at shell closure along Ni isotopic chain Malbrunot-Ettenauer et al. (2022). As demonstrated in Ref. Sharma et al. (1995), shell closure effect results from the rather small isospin dependent of spin-orbit interactions. Meanwhile, the great deviation are also observed in charge radii at the proton number $Z=26$ and $28$. For WS model, the shell effect and isospin effect in nuclear charge radii are systematically investigated Wang and Li (2013). Along $N=32$ and $34$ isotonic chains, the $Z=20$ and $28$ shell closure are observed apparently. Furthermore, the rapid increase of charge radii can also be found across $Z=20$, but weakened in our calculations. And the charge radii of ${}^{51}$ and ${}^{52}$a isotopes are underestimated in the WS model. From $Z=29$-$31$ isotopic chains, charge radii of $N=32$ isotones can be reproduced well by NL3, PK1, $*{}^{*}$, BNN+D4 and WS models. But the results obtained by HFB-21 are slightly overestimated against the experimental data.
@@
In general, the odd-even staggering in charge radii describes the fact that the nuclear charge radii of odd-$N$ isotopes are smaller than the averages of their even-$N$ neighbours. In order to emphasize these local variations, a three-point formula has been employed to extract the odd-even oscillation behaviors of nuclear charge radii along a specific isotopic chain Reinhard and Nazarewicz (2017); An et al. (2020). As shown in Fig. 3, the abrupt changes of charge radii across $Z=22$ obtained by NL3, PK1 and $*{}^{*}$ effective forces are also shown along $N=32$ and $34$ isotones, but later with a less slope. For further inspecting this sudden change in nuclear charge radii, the OES of charge radii along $N=28$, $32$ and $34$ isotones are also calculated by this three-point formula. The definition is rewritten as $\Delta_{r}(Z,N)=\frac{1}{2}[2R(Z,N)-R(Z-1,N)-R(Z+1,N)]$, where $R(Z,N)$ is rms charge radius along isotonic chain.
@@
As shown in Fig. 4, the OES of charge radii of $N=28$, $32$ and $34$ isotones as a function of proton number are plotted within NL3, PK1 and $*{}^{*}$ effective forces. For the $N=28$ isotones, the results obtained by NL3 and $*{}^{*}$ forces against well the experimental data along $Z=22$-$24$, while PK1 force give the opposite sign until $Z=25$. At $Z=20$ and $28$, these three methods give the similar trend, in which the OES of charge radii presents the anomalous behavior. As demonstrated in Ref. An et al. (2022b), this results from the strong shell closure effect.
@@
Along $N=32$ and $34$ isotones, NL3, PK1 and $*{}^{*}$ effective forces can reproduce the experimental data from $Z=29$ to $30$. Both the NL3 and $*{}^{*}$ forces give the almost identical results, while PK1 give the opposite sign at $Z=27$ along $N=32$ isotone. The same scenarios can also be found at $Z=24$ along $N=34$ isotone. Here, we greatly pay more attention to the local variations of charge radii across the proton number $Z=22$. For $N=32$ and $34$ isotones, one can find that the amplitudes of OES in charge radii are increased at $Z=22$. This is consistent with the trend of changes of charge radii as shown in Fig. 3. In addition, one can find that the amplitudes of OES in charge radii are comparative for $Z=20$ and $Z=22$ isotones.

### v2

@@ III.1 Shell quenching of nuclear charge radii at $N=28$
where $R(N,Z)$ is rms charge radius for a nucleus with neutron number $N$ and proton number $Z$. In Fig. 1 1(c) and 1(d), the OES of charge radii in potassium and calcium isotopes is depicted. For potassium isotopes, the OES in charge radii can be reproduced well across $N=20$ shell closure. Extending to neutron-deficient regions, the obtained OES amplitudes for these three effective forces cannot reproduce the experimental data well. This can be due to the fact that the charge radii in neutron-deficient region are overestimated by NL3 and NL3 sets and underestimated by PK1 force. For neutron-rich region, the PK1 force gives the enlarged OES of charge radii with respect to NL3 and NL3 sets. Along calcium isotopes, the OES behaviors of charge radii for 37-47Ca isotopes can be reproduced well by these forces. Across the $N=28$ filled shell, PK1 set gives the inverse OES amplitudes in charge radii with respect to NL3 and NL3 sets. Besides, it can be found that the OES behavior is weakened at the neutron number $N=28$. This is in accord with Ref. An et al. (2022b) where the weakened OES in charge radii can be regarded as signature of the emergence of neutron magic numbers.
@@
Figure 2: Same as Fig. 1, but for (a) scandium, (b)titanium, (c) vanadium, (d) chromium, (e) manganese, (f) iron, (g) cobalt, (h) nickel, and (i) copper isotopes. Experimental data are taken from Refs. Angeli and Marinova (2013); Li et al. (2021), while the data of Sc isotopes are taken from Ref. König et al. (2023), the value of 54Ni is extracted from Ref. Pineda et al. (2021) and 58-68,70Ni are extracted from Ref. Malbrunot-Ettenauer et al. (2022).
@@
To facilitate the instructive understanding of charge radii on the isotopic chains of neighboring elements, the systematic evolution of charge radii along $Z=21$-$29$ isotopic chains is plotted in Fig. 2. From this figure, one can find that the shell effect at $N=28$ is evidently observed. Across $N=28$, the abrupt changes are universally observed in this region, while the slope of change is lower for scandium ($Z=21$) and titanium ($Z=22$). In Fig. 2 (a), these three forces give the inverted parabolic-like behaviors in charge radii between $N=20$ and $N=28$, while PK1 parameter set slightly underestimates the results. The latest study suggests that the charge radii of 41Sc isotope is surprisingly shrunk König et al. (2023), then the abrupt increase is occurred beyond the neutron number $N=20$. In our calculations, the PK1 force can reproduce the charge radii of 40,41Sc isotopes, but the values of 42-46Sc isotopes are exaggeratedly underestimated. By contrast, these phenomena are mostly met by the NL3 and NL3 forces. The NL3 and NL3 forces can cover the experimental uncertainties of 40,41Sc, but the values of 42,43Sc are underestimated. It is notably mentioned that our calculations can give the trend of kink, but the amplitude is heavily reduced. Besides, the last unpaired nucleon is just tackled by the blocking approximation, namely occupies the last single-particle level. This may also result in the underestimated results. The same scenarios are also encountered for titanium isotopes. As shown in Ref. Gangrsky et al. (2004), the data of the charge radii on the titanium isotopes show a continuous decrease from $N=22$ to $28$. The calculated results for 44-46Ti are slightly deviated from experimental data. But the trend of changes of charge radii are not changed, especially for 44-46Ti where the odd-even oscillation behaviors can be reproduced well.
@@
For V, Cr, and Co isotopic chains, the experimental data can be reproduced well. However, more experimental data are urgently encouraged to support our results. In these regions the overall slope increases gradually with the increasing neutron numbers. Also the shell closure effect in charge radii is clearly pronounced at $N=28$ due to the rather small isospin dependence of spin-orbit interactions Sharma et al. (1995). Along manganese isotopes, results obtained by NL3, PK1, and NL3 forces can reproduce the experimental data across $N=28$ shell closure. However, there exists a slight deviation between experimental data and results obtained by PK1 force. For 50,51Mn and 52Fe, the calculated results are all reduced. Meanwhile, the data of 56-58Fe isotopes are also slightly underestimated.
@@
Recent studies on nuclear size are performed along nickel isotopic chain Pineda et al. (2021); Malbrunot-Ettenauer et al. (2022). In Fig. 2 (h), one can find that NL3, PK1, and NL3 forces can describe the general trend of changes of charge radii along nickel isotopic chain. Especially the slight odd-even oscillation behavior can be reproduced well against experimental data. In Ref. Malbrunot-Ettenauer et al. (2022), it is mentioned that the trend of changes of charge radii for ${}^{62-70}\mathrm{Ni}$ isotopes are overestimated by Fayans EDF model. The sophisticated Fayans energy density functional (EDF) can reproduce the staggering effects of charge radii for Ca isotopes Reinhard and Nazarewicz (2017). Furthermore, the inverted parabolic-like behavior between $N=20$ and $N=28$ can also be reproduced well. As demonstrated in Ref. Malbrunot-Ettenauer et al. (2022), the isovector component is excluded in pairing interactions. This may lead to the lack of an isovector component in its pairing part. In Ref. Kortelainen et al. (2022), the charge radii of even-even Ca-Zn nuclei are obtained by the $ab$ $initio$ coupled cluster theory and nuclear density functional theory, in which a universal pattern from neutron numbers $N=28$ to $N=40$ is exhibited, namely the change in charge radii shows weak dependence on the atomic number. Actually, our results are in accord with those.
@@
The pioneering work shows that the charge radii formula without considering the unpaired nucleons correlation gives the overestimated OES of charge radii in copper isotopes An et al. (2022c). As shown in Fig. 2 (i), the improved formula can reduce the overestimated odd-even oscillations of charge radii. In Ref. de Groote et al. (2020), it is found that the charge radii along copper isotopic chain are also slightly overestimated by Fayans DFT model. Especially close to neutron-rich region, these overestimated OES behaviors are evidently represented. Therefore, the further performance of optimization-based machine learning method is addressed through Fayans EDF model Bollapragada et al. (2021).
III.2 Charge radii of $N=28$, $32$, and $34$ isotones
@@
The emergence of rapidly increasing charge radii is commonly observed across the traditional neutron magic numbers $N=28$, $50$, $82$, and $126$ in the nuclear chart Kreim et al. (2014); Gorges et al. (2019); Goddard et al. (2013); Procter et al. (2012). These phenomena arise from the relative stable binding properties with respect to the adjacent nuclei. Meanwhile, the new magicity of $N=32$ and $34$ has also been identified by multi-aspects. Recent studies show that the isotopes with neutron number $N=32$ are proposed to exhibit localized magic behavior due to an observed sudden decrease in binding energy beyond $N=32$, such as in 51K ($Z=19$Rosenbusch et al. (2015), 52Ca ($Z=20$An et al. (2018); Wienholtz et al. (2013); Gallant et al. (2012), 53Sc ($Z=21$Xu et al. (2019), and 54Ti ($Z=22$Janssens et al. (2002); Leistenschneider et al. (2018). This new magicity is also identified from the high excitation energy of the first excited state Steppenbeck et al. (2015); Huck et al. (1985); Liddick et al. (2004); Prisciandaro et al. (2001) and the reduced $B(E2)$ transition probabilities Goldkuhle et al. (2019); Seidlitz et al. (2011). Especially for 52Ca, the doubly magic nature is also confirmed by two-proton knockout reaction Gade et al. (2006). As demonstrated in Ref. Reiter et al. (2018), the signature of $N=32$ sub-shell closure is not identified across the vanadium isotopic chain ($Z=23$) and higher proton numbers.
@@
Figure 3: Charge radii of $N=28$ (open circle), $32$ (open triangle) and $34$ (open square) isotones as a function of proton number are plotted within (a) NL3, (b) PK1, and (c) NL3 forces. The results obtained by the (d) BNN+D4 Dong et al. (2022), (e) HFB-21 Goriely et al. (2010), and (f) WS Wang and Li (2013) models are also shown for comparison. The corresponding experimental data are taken from Refs. Angeli and Marinova (2013); Li et al. (2021) (solid marks).
@@
In Fig. 3, the charge radii of nuclei along $N=28$, $32$, and $34$ isotonic chains are plotted. It clearly shows that the charge radii monotonically increase with the increasing proton numbers. As the slope of these curves involving different proton numbers is much changed, i.e., the local variations can be found obviously. The charge radii almost changes linearly from K ($Z=19$) to Fe ($Z=26$) along $N=28$ isotonic chain. The shrinking trend occurs until nickel ($Z=28$), and then the abrupt change occurs across $Z=28$. This result suggests that the shell closure effect of charge radii can also be found along $N=28$ isotones. These calculations for the rms charge radii reveal the characteristic kink at the $Z=28$ shell closure in accordance with the corresponding experimental radii.
@@
Along $N=32$ isotonic chain, the abrupt changes of charge radii across nuclei with $Z=22$ are observed within NL3 [Fig. 3 (a)], PK1 [Fig. 3(b)], and NL3 [Fig. 3(c)] effective forces. From V to Ga isotopes, the charge radii are changed linearly with the increasing proton numbers, but the results of 58Fe are slightly deviated. This can be seen from the Fig. 2 that charge radii of Fe isotopes are systematically underestimated by these three forces. From K to Ti isotopes, the increasing trend of charge radii are almost linear. Same scenario can also be found along $N=34$ isotonic chain, but the slope of changes of charge radii from $Z=22$ to $Z=23$ is smaller. As discussed above, magicity means the nuclei process relatively stable properties. This pronounced feature can also be remarkably reflected by nuclear size. As suggested in Ref. Reiter et al. (2018), the signature of $N=32$ sub-shell closure is not identified across $Z=23$ and higher proton numbers. This is consistent with our results where the charge radii keep linear increase after the abrupt increase at $Z=22$. This may provide a signature to verify the new magicity of nuclei with $N=32$. The same scenario can also be observed along $N=34$ isotonic chain, but the slope of changes of charge radii is smaller. Consequently, more accurate experimental data are urgently needed in proceeding future.
@@
The latest experimental data show that there has been no sharp change in charge radii across $N=32$ Koszorús et al. (2021). Although the magicity of nuclei with $N=32$ has been performed, the shrinking phenomena are not observed along potassium Koszorús et al. (2019) and calcium Garcia Ruiz et al. (2016) isotopes. A more complete picture may shed light on the unexplained phenomena and reveal details of the proton-neutron interaction under the influence of the $N=32$ and $34$ sub-shell closure. The $N=34$ sub-shell closure has been so far performed in Ca Michimasa et al. (2018); Steppenbeck et al. (2013); Chen et al. (2019) and Ar Liu et al. (2019) isotopes. However, the slightly sharp transition in charge radii is also observed across $Z=22$. A quantitative theoretical interpretation of the well-known sub-shell effect in $N$ has been hampered by the fact that detailed calculations use somewhat inconsistent models in treating even-$N$ and odd-$N$ nuclei. The present case of a pronounced $Z$ dependence of the charge radii along $N=28$, $32$, and $34$ isotonic chains may provide the clue for a clear-cut phenomenological interpretation.
@@
Figure 4: Odd-even staggering of charge radii versus (a) $N=28$, (b) $32$, and (c) $34$ isotones as a function of proton number are plotted within NL3 (open diamond), PK1 (open square), and NL3 (solid triangle) forces. The corresponding experimental data are taken from Refs. Angeli and Marinova (2013); Li et al. (2021).
@@
Such a comparison should be helpful in providing an accessible understanding of nuclear charge radii and in making reliable predictions for those nuclei lacking an experimental value. Although the systematic evolution of the charge radii over a large range of proton numbers seems to be consistent with the presence of a noticeable sub-shell effects at $N=32$ and $34$, it misses the pronounced experimental data to identify these results. In order to further examine the universal trend of nuclear charge radii, it will be also instructive to gather similar data on the isotonic chains at neutron number $N=28$, $32$, and $34$. In Figs. 3(d), 3(e), and 3(f), the results obtained by Bayesian neural network (BNN+D4) Dong et al. (2022), HFB-21 Goriely et al. (2010), and WS Wang and Li (2013) are also shown for comparison.
@@
In Fig. 3 (d), the change of charge radii obtained by BNN+D4 is almost linear along $N=28$ isotonic chain. The abrupt increases are predicted along $N=32$ and $34$ isotonic chains at the proton numbers $Z=20$ and $28$, respectively. This can be understood clearly where the pronounced shell effects are emphasized in the calibrated protocol Dong et al. (2022). For HFB-21 model Goriely et al. (2010), the charge radii of $N=28$, $32$, and $34$ isotonic chains change linearly with the increasing proton numbers. While the slight shrinking occurs at $Z=27$, this may be a result from the unpaired proton. This means the strong shell effect is weakened in Skyrme functionals. As demonstrated in Refs. Tajima et al. (1993); Gorges et al. (2019), the calculations in non-relativistic density functional theories (DFT) based on conventional functionals are unable to reproduce the kink well at magic numbers. Such as Skyrme functionals utilizing SV-min force generally cannot reproduce the characteristic kink typically found at shell closure along Ni isotopic chain Malbrunot-Ettenauer et al. (2022). As demonstrated in Ref. Sharma et al. (1995), shell closure effect results from the rather small isospin dependence of spin-orbit interactions. Meanwhile, the great deviations are also observed in charge radii at the proton numbers $Z=26$ and $28$. For WS model, the shell effect and isospin effect in nuclear charge radii are systematically investigated Wang and Li (2013). Along $N=32$ and $34$ isotonic chains, the $Z=20$ and $28$ shell closure are observed apparently. Furthermore, the rapid increase of charge radii can also be found across $Z=20$, but weakened in our calculations. And the charge radii of 51K and 52Ca isotopes are underestimated in the WS model. From $Z=29$-$31$ isotopic chains, charge radii of $N=32$ isotones can be reproduced well by NL3, PK1, NL3, BNN+D4, and WS models. But the results obtained by HFB-21 are slightly overestimated against the experimental data.
Considering the neutron-proton correlation in describing the evolution of nuclear charge radii, the increasing trends of charge radii along $N=32$ and $34$ isotonic chains seem to suggest the limitation of new magicity at the proton number $Z=22$. Meanwhile, it leads to the vanished magicity of $Z=28$ along the neutron numbers $N=32$ and $N=34$ isotones. It is also mentioned that the magicity at $Z=28$ is weakened along $N=32$ and $34$ isotonic chains for the HFB-21 model, but the BNN+D4 and WS models show the magicity at $Z=28$ as well as at $Z=20$.
@@
In general, the odd-even staggering in charge radii describes the fact that the nuclear charge radii of odd-$N$ isotopes are smaller than the averages of their even-$N$ neighbours. In order to emphasize these local variations, a three-point formula has been employed to extract the odd-even oscillation behaviors of nuclear charge radii along a specific isotopic chain Reinhard and Nazarewicz (2017); An et al. (2020). As shown in Fig. 3, the abrupt changes of charge radii across $Z=22$ obtained by NL3, PK1, and NL3 effective forces are also shown along $N=32$ and $34$ isotones, but the latter with a less slope. For further inspecting this sudden change in nuclear charge radii, the OES of charge radii along $N=28$, $32$, and $34$ isotones are also calculated by this three-point formula. The definition is rewritten as $\Delta_{r}(Z,N)=\frac{1}{2}[2R(Z,N)-R(Z-1,N)-R(Z+1,N)]$, where $R(Z,N)$ is rms charge radius along isotonic chain.
@@
As shown in Fig. 4, the OES of charge radii of $N=28$, $32$, and $34$ isotones as a function of proton number are plotted within NL3, PK1, and NL3 effective forces. For the $N=28$ isotones,the results obtained by NL3 and NL3 forces agree well against the experimental data along $Z=23$-$25$ isotopes, while PK1 force give the opposite sign until $Z=26$. At $Z=20$ and $28$, these three methods gives the similar trend, in which the OES of charge radii presents the anomalous behavior. As demonstrated in Ref. An et al. (2022b), this results from the strong shell closure effect.
@@
Along $N=32$ and $34$ isotones, NL3, PK1, and NL3 effective forces can reproduce the experimental data from $Z=29$-$30$. Both the NL3 and NL3 forces give the almost identical results, while PK1 gives the opposite sign at $Z=27$ along $N=32$ isotone. The same scenarios can also be found at $Z=24$ along $N=34$ isotone. Here, we greatly pay more attention to the local variations of charge radii across the proton number $Z=22$. For $N=32$ and $34$ isotones, one can find that the amplitudes of OES in charge radii are increased at $Z=22$. This is consistent with the trend of changes of charge radii as shown in Fig. 3. In addition, one can find that the amplitudes of OES in charge radii are comparative for $Z=20$ and $Z=22$ isotones.